# Leveraging Personalized PageRank and Higher-Order Topological Structures for Heterophily Mitigation in Graph Neural Networks

### v1

@@ 3.1 Background
Consider an undirected, unweighted graph $G=(V,E)$, where $V=\{v_{1},v_{2},\dots,v_{N}\}$ is a finite set of $N=|V|$ nodes, and $E\subseteq V\times V$ is the set of $M=|E|$ edges. Let $X\in\mathbb{R}^{N\times F}$ denote the node feature matrix, where $F$ is the feature dimension, and $A\in\{0,1\}^{N\times N}$ represent the adjacency matrix, with $A_{ij}=1$ if $(v_{i},v_{j})\in E$ and 00 otherwise.
@@ 4.1 Constructing Higher-order Network
To capture higher-order interactions, we transform a graph $G=(V,E)$ into simplicial complexes(SCs) via clique complex lifting Schaub et al. (2020). Let $C(G)$ denote the set of all cliques in $G$, where a clique $C_{i}\subseteq V$ is a complete subgraph. Each $p$-clique in $C(G)$ is mapped to a $(p-1)$-simplex, where 2-simplices are the 3-cliques in the graph, forming the SC $\mathcal{K}=\bigcup_{C_{i}\in C(G)}\sigma_{i}$. This process encodes multi-node interactions beyond pairwise edges.
@@ 4.2.1 Higher-order Information Integration in PPR
where $I^{p}$ is the identity matrix corresponding to the number of simplices of order $p$ and $\tilde{A}_{p}$ is the normalized adjacency matrix incorporating the higher-order simplicial structure, representing the adjacency matrix between 00-simplices(nodes) and $p$-simplices (where $p>0$) . This innovative formula extends the basic idea of PPR to account for higher-order relationships within the simplicial complex.
@@ 4.3.2 Incorporating Higher-order Structures
[…]
@@ 4.3.3 Higher-order Spectral Convolution Formulation
[…]

### v2

@@ 3.1 Background
Consider an undirected, unweighted graph $G=(V,E)$, where $V=\{v_{1},v_{2},\dots,v_{N}\}$ is a finite set of $N=|V|$ nodes, and $E\subseteq V\times V$ is the set of $M=|E|$ edges. Let $X\in\mathbb{R}^{N\times F}$ denote the node feature matrix, where $F$ is the feature dimension, and $A\in\{0,1\}^{N\times N}$ represent the adjacency matrix, with $A_{ij}=1$ if $(v_{i},v_{j})\in E$ and 0 otherwise.
@@ 4.1 Constructing Higher-order Network
To capture higher-order interactions, we transform a graph $G=(V,E)$ into simplicial complexes(SCs) via clique complex lifting Schaub et al. (2020). Let $C(G)$ denote the set of all cliques in $G$, where a clique $C_{i}\subseteq V$ is a complete subgraph. Each $p$-clique in $C(G)$ is mapped to a (p1)(p-1)-simplex, where 2-simplices are the 3-cliques in the graph, forming the SC $\mathcal{K}=\bigcup_{C_{i}\in C(G)}\sigma_{i}$. This process encodes multi-node interactions beyond pairwise edges.
@@ 4.2.1 Higher-order Information Integration in PPR
where $I^{p}$ is the identity matrix corresponding to the number of simplices of order $p$ and $\tilde{A}_{p}$ is the normalized adjacency matrix incorporating the higher-order simplicial structure, representing the adjacency matrix between 0-simplices(nodes) and $p$-simplices (where $p>0$) . This innovative formula extends the basic idea of PPR to account for higher-order relationships within the simplicial complex.
@@ 4.3.2 Incorporating Higher-order Structures
where $Z^{p}$ represents the output combining contributions from the $k$-th hop, $\beta_{k}$ are $k$ learnable coefficients, $\tilde{\Pi}^{p}_{HiPPR}$ captures neighborhood information at order $p$, $\tilde{\Pi}^{p}_{HiPPR}$ is the normalised version of ${\Pi}^{p}_{HiPPR}$, formally defined as $\tilde{\Pi}^{p}_{HiPPR}=D^{-1/2}{\Pi}^{p}_{HiPPR}D^{-1/2}$, where $D$ is the diagonal degree matrix. $X$ is the input feature matrix, $\rho$ is the simplified linear functions suggested byWang and Zhang (2022), and $\big\|$ denotes concatenation.
@@ 4.3.3 Higher-order Spectral Convolution Formulation
where $\beta_{p,k}$, $\Theta_{p}$ and $W$ are learnable parameters, $\big\|$ denotes concatenation, and $P$ refers to the highest order of the simplicial complex under consideration.